# Learning 2D Invariant Affordance Knowledge for 3D Affordance Grounding

### v2

@@ 2.1 Affordance Learning
Affordance learning is crucial in robotics, particularly for manipulating articulated objects (Ning et al. 2024). Consequently, many studies focus on detecting the affordance areas of target objects or scenes (Fang et al. 2018; Luo et al. 2022a, b; Nagarajan, Feichtenhofer, and Grauman 2019). Some works extract affordance knowledge from 2D data, i.e., images and videos (Chuang et al. 2018; Do, Nguyen, and Reid 2018; Li et al. 2023; Luo et al. 2021; Roy and Todorovic 2016; Thermos, Daras, and Potamianos 2020; Zhao, Cao, and Kang 2020; Chen et al. 2023; Mur-Labadia, Guerrero, and Martinez-Cantin 2023), while others explore affordance using natural language guidance (Lu et al. 2022; Mi et al. 2020, 2019; Chen, Cong, and Kan 2024; Yoshida et al. 2024). These methods aim to predict the affordance regions of 2D targets. However, robotics tasks often require 3D information, and transferring affordance knowledge learned from 2D data to 3D scenarios can lead to failures in real-world applications. As a result, several 3D affordance-related datasets have been proposed (Geng et al. 2023b; Liu et al. 2022; Mo et al. 2019), and some studies focus on leveraging these 3D datasets to ground object affordance (Mo et al. 2022; Nagarajan and Grauman 2020; Xu et al. 2022a; Delitzas et al. 2024). 3D AffordanceNet (Deng et al. 2021) first introduces a benchmark dataset for grounding affordance regions on object point clouds. Building on this 3D dataset, IAGNet (Yang et al. 2023b) proposes a method for learning 3D affordance from reference images. However, the significant appearance variations among reference images and 3D point clouds pose a challenge. LASO (Li et al. 2024b) deals with this by replacing images with natural language descriptions of the interaction area. However, completely discarding the vision information may overlook distinct affordance area guidance. Therefore, we propose a method that better aligns visual latent features from multiple images to effectively guide affordance grounding. Where2Explore (Ning et al. 2024) acquires affordance information by explicitly estimating the geometric similarity across different categories, thereby enhancing generalization. GAPartNet (Geng et al. 2023b) introduces cross-category tasks and a dataset to explore the consistency of generalizable and actionable parts. PartManip (Geng et al. 2023a) identifies actionable parts to facilitate cross-category object manipulation. While these approaches do not fully capture the category-level consistency features crucial for affordance grounding. By disregarding inconsistent parts across images and implicitly emphasizing consistency, our method more effectively learns affordance information from reference images.
@@ 2.2 Image-Point Cloud Cross-Modal Learning
Cross-modal tasks enhance individual modalities by incorporating information from one or more other modalities. These additional modalities can serve as conditions to guide the learning progress, often resulting in a positive impact (Qu et al. 2024). The combination of images and point clouds is particularly complementary in downstream tasks: point clouds provide stereospatial information, while images offer rich color and texture details. To leverage this complementarity, many works focus on aligning each pixel to a corresponding point to enrich the semantic information of raw data (Zhuang et al. 2021; Zhao et al. 2021; Xu et al. 2022b; Vora et al. 2020; Tan et al. 2021; Krispel et al. 2020; Chen et al. 2024b). Additionally, some methods aim to learn multi-view fusion with point clouds, seeking common knowledge and aligning both image-to-image and image-to-point cloud (Jaritz, Gu, and Su 2019; Zhao, Lu, and Zhou 2021; Li et al. 2022; Yang et al. 2023a; Wang et al. 2024; Jiang et al. 2024; Chen et al. 2024a; Zhang et al. 2024a; Xu et al. 2024; Jiao et al. 2024). However, unlike these approaches, which may not fully capture the relationships between multiple images and often lack sufficient modeling of consistent image features for effective fusion with point clouds, our method first establishes consistency between images and then aligns the invariant image features with the point cloud, leading to a more robust integration of visual and spatial information.
@@ 3.2 Invariant Affordance Knowledge Extraction Module
where $\mathbf{Q}^{(l-1)}_{i}\in\mathbb{R}^{M\times C}$ and $\mathbf{F}^{(l-1)}_{i}\in\mathbb{R}^{D\times H\times W}$ denote the affordance queries and reference image features from layer $(l-1)$, respectively. Then, to leverage the inherent consistency among multiple affordance queries, we apply an $\operatorname{MLP}$ layer to align and aggregate all the queries. The aggregated features are then fed back into the image feature branch by multi-head cross attention. Additionally, to continually extract consistent image features, the image feature branch is also updated iteratively using a multi-head self-attention layer. This process can be expressed as follows:
@@ 3.4 Decoder and Loss Function
where $\mathbf{P}_{pred}\in\mathbb{R}^{N\times 1}$. The total loss consists of three components: $\mathbf{\mathcal{L}}_{CE}$, $\mathbf{\mathcal{L}}_{Sim}$ and $\mathbf{\mathcal{L}}_{HM}$. $\mathbf{\mathcal{L}}_{CE}$ represents the cross-entropy loss between $y$ and $\hat{y}$. $\mathbf{\mathcal{L}}_{Sim}$ is the cosine similarity loss calculated at each layer of $\{\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{n}\}$, which aims to align image features while obtaining invariant affordance knowledge. $\mathbf{\mathcal{L}}_{HM}$ combines focal loss (Lin et al. 2017) with dice loss (Milletari, Navab, and Ahmadi 2016), and is calculated between $\mathbf{P}_{pred}$ and $\mathbf{P}_{label}$, supervising the point-wise heatmap on point clouds. The total loss is formulated as follows:
@@ 4.1 Experimental Settings
The most relevant works are IAGNet (Yang et al. 2023b) and LASO (Li et al. 2024b), both of which leverage 3D AffordanceNet (Deng et al. 2021) to derive the corresponding 3D shapes. In addition, consistent with IAGNet, we adopt SOTA image-3D cross-modal methods as baselines for a more comprehensive evaluation, encompassing PMF (Zhuang et al. 2021), XMF (Aiello, Valsesia, and Magli 2022), and ILN (Chen et al. 2022), which concentrate on fusing image and point features, as well as MBDF (Tan et al. 2021), FRCNN (Xu et al. 2022b), and PFusion (Xu, Anguelov, and Jain 2018) which are predominantly employed for multi-modal object grounding. We reproduce these methods with the same feature extractors and settings in the original papers to ensure a fair comparison. Besides, we enhance these methods by replacing the single image or text inputs with multiple ones, allowing for a more direct comparison with our dataset. All the evaluation metrics follow previous works: Area Under the Curve (AUC(Lobo, Jiménez-Valverde, and Real 2008), average Intersection Over Union (aIOU(Rahman and Wang 2016), SIMilarity (SIM(Swain and Ballard 1991) and Mean Absolute Error (MAE(Willmott and Matsuura 2005).
@@ 4.2 Quantitative Analysis
Table 1 reports the metrics of the proposed method MIFAG compared with other methods on the MIPA benchmark. The results demonstrate that the proposed MIFAG achieves state-of-the-art performance, significantly surpassing existing methods in 3D point cloud affordance prediction. Specifically, MIFAG achieve scores of 85.10 in AUC and 20.50 in aIOU, outperforming the second-best method, LASO (Li et al. 2024b), and IAGNet (Yang et al. 2023b), with an improvement of +1.97 in AUC and +2.58 in aIOU. Notably, our method also demonstrates impressive affordance capabilities in unseen point cloud prediction, outperforming the second-best method, IAGNet by 1.22 in AUC and LASO by 0.05 in aIOU. Additionally, 3D segmentation and tracking methods such as PMF (Zhuang et al. 2021), XMF (Aiello, Valsesia, and Magli 2022), and MBDF (Tan et al. 2021) perform significantly worse compared to the affordance method.
@@ B.3 Loss Functions
Consistent with IAGNet (Yang et al. 2023b), the heatmap loss $\mathcal{L}_{HM}$ consists of a focal loss (Lin et al. 2017) and a dice loss (Milletari, Navab, and Ahmadi 2016), which are used to supervise the predicted affordance heatmap and the corresponding ground truth. This can be formulated as follows:
@@ B.4 Evaluation Metrics
Here, we provide a detailed explanation of the evaluation metrics used in our study. Following previous research, we adopt AUC (Lobo, Jiménez-Valverde, and Real 2008), aIOU (Rahman and Wang 2016), SIM (Swain and Ballard 1991), and MAE (Willmott and Matsuura 2005) as our evaluation metrics:
@@
    IOU=Intersction AreaUnion $\text{IOU}=\frac{\text{Intersction Area}}{\text{Union Area}}.$ (21)

### v3

@@ 2.1 Affordance Learning
Affordance learning is crucial in robotics, particularly for manipulating articulated objects (Ning et al. 2024). Consequently, many studies focus on detecting the affordance areas of target objects or scenes (Fang et al. 2018; Luo et al. 2022a, b; Nagarajan et al. 2019). Some works extract affordance knowledge from 2D data, i.e., images and videos (Chuang et al. 2018; Do et al. 2018; Li et al. 2023; Luo et al. 2021; Roy and Todorovic 2016; Thermos et al. 2020; Zhao et al. 2020; Chen et al. 2023; Mur-Labadia et al. 2023), while others explore affordance using natural language guidance (Lu et al. 2022; Mi et al. 2020, 2019; Chen et al. 2024a; Yoshida et al. 2024). These methods aim to predict the affordance regions of 2D targets. However, robotics tasks often require 3D information, and transferring affordance knowledge learned from 2D data to 3D scenarios can lead to failures in real-world applications. As a result, several 3D affordance-related datasets have been proposed (Geng et al. 2023b; Liu et al. 2022; Mo et al. 2019), and some studies focus on leveraging these 3D datasets to ground object affordance (Mo et al. 2022; Nagarajan and Grauman 2020; Xu et al. 2022a; Delitzas et al. 2024). 3D AffordanceNet (Deng et al. 2021) first introduces a benchmark dataset for grounding affordance regions on object point clouds. Building on this 3D dataset, IAGNet (Yang et al. 2023b) proposes a method for learning 3D affordance from reference images. However, the significant appearance variations among reference images and 3D point clouds pose a challenge. LASO (Li et al. 2024b) deals with this by replacing images with natural language descriptions of the interaction area. However, completely discarding the vision information may overlook distinct affordance area guidance. Therefore, we propose a method that better aligns visual latent features from multiple images to effectively guide affordance grounding. Where2Explore (Ning et al. 2024) acquires affordance information by explicitly estimating the geometric similarity across different categories, thereby enhancing generalization. GAPartNet (Geng et al. 2023b) introduces cross-category tasks and a dataset to explore the consistency of generalizable and actionable parts. PartManip (Geng et al. 2023a) identifies actionable parts to facilitate cross-category object manipulation. While these approaches do not fully capture the category-level consistency features crucial for affordance grounding. By disregarding inconsistent parts across images and implicitly emphasizing consistency, our method more effectively learns affordance information from reference images.
@@ 2.2 Image-Point Cloud Cross-Modal Learning
Cross-modal tasks enhance individual modalities by incorporating information from one or more other modalities. These additional modalities can serve as conditions to guide the learning progress, often resulting in a positive impact (Qu et al. 2024). The combination of images and point clouds is particularly complementary in downstream tasks: point clouds provide stereospatial information, while images offer rich color and texture details. To leverage this complementarity, many works focus on aligning each pixel to a corresponding point to enrich the semantic information of raw data (Zhuang et al. 2021; Zhao et al. 2021a; Xu et al. 2022b; Vora et al. 2020; Tan et al. 2021; Krispel et al. 2020; Chen et al. 2024c). Additionally, some methods aim to learn multi-view fusion with point clouds, seeking common knowledge and aligning both image-to-image and image-to-point cloud (Jaritz et al. 2019; Zhao et al. 2021b; Li et al. 2022; Yang et al. 2023a; Wang et al. 2024; Jiang et al. 2024; Chen et al. 2024b; Zhang et al. 2024a; Xu et al. 2024; Jiao et al. 2024). However, unlike these approaches, which may not fully capture the relationships between multiple images and often lack sufficient modeling of consistent image features for effective fusion with point clouds, our method first establishes consistency between images and then aligns the invariant image features with the point cloud, leading to a more robust integration of visual and spatial information.
@@ 3.2 Invariant Affordance Knowledge Extraction Module
where $\mathbf{Q}^{(l-1)}_{i}\in\mathbb{R}^{M\times C}$ and $\mathbf{F}^{(l-1)}_{i}\in\mathbb{R}^{D\times H\times W}$ denote the affordance queries and reference image features from layer (l1)(l-1), respectively. Then, to leverage the inherent consistency among multiple affordance queries, we apply an $\operatorname{MLP}$ layer to align and aggregate all the queries. The aggregated features are then fed back into the image feature branch by multi-head cross attention. Additionally, to continually extract consistent image features, the image feature branch is also updated iteratively using a multi-head self-attention layer. This process can be expressed as follows:
@@ 3.4 Decoder and Loss Function
where $\mathbf{P}_{pred}\in\mathbb{R}^{N\times 1}$. The total loss consists of three components: $\mathbf{\mathcal{L}}_{CE}$, $\mathbf{\mathcal{L}}_{Sim}$ and $\mathbf{\mathcal{L}}_{HM}$. $\mathbf{\mathcal{L}}_{CE}$ represents the cross-entropy loss between $y$ and $\hat{y}$. $\mathbf{\mathcal{L}}_{Sim}$ is the cosine similarity loss calculated at each layer of $\{\mathbf{F}_{1},\mathbf{F}_{2},\dots,\mathbf{F}_{n}\}$, which aims to align image features while obtaining invariant affordance knowledge. $\mathbf{\mathcal{L}}_{HM}$ combines focal loss (Lin et al. 2017) with dice loss (Milletari et al. 2016), and is calculated between $\mathbf{P}_{pred}$ and $\mathbf{P}_{label}$, supervising the point-wise heatmap on point clouds. The total loss is formulated as follows:
@@ 4.1 Experimental Settings
The most relevant works are IAGNet (Yang et al. 2023b) and LASO (Li et al. 2024b), both of which leverage 3D AffordanceNet (Deng et al. 2021) to derive the corresponding 3D shapes. In addition, consistent with IAGNet, we adopt SOTA image-3D cross-modal methods as baselines for a more comprehensive evaluation, encompassing PMF (Zhuang et al. 2021), XMF (Aiello et al. 2022), and ILN (Chen et al. 2022), which concentrate on fusing image and point features, as well as MBDF (Tan et al. 2021), FRCNN (Xu et al. 2022b), and PFusion (Xu et al. 2018) which are predominantly employed for multi-modal object grounding. We reproduce these methods with the same feature extractors and settings in the original papers to ensure a fair comparison. Besides, we enhance these methods by replacing the single image or text inputs with multiple ones, allowing for a more direct comparison with our dataset. All the evaluation metrics follow previous works: Area Under the Curve (AUC(Lobo et al. 2008), average Intersection Over Union (aIOU(Rahman and Wang 2016), SIMilarity (SIM(Swain and Ballard 1991) and Mean Absolute Error (MAE(Willmott and Matsuura 2005).
@@ 4.2 Quantitative Analysis
Table 1 reports the metrics of the proposed method MIFAG compared with other methods on the MIPA benchmark. The results demonstrate that the proposed MIFAG achieves state-of-the-art performance, significantly surpassing existing methods in 3D point cloud affordance prediction. Specifically, MIFAG achieve scores of 85.10 in AUC and 20.50 in aIOU, outperforming the second-best method, LASO (Li et al. 2024b), and IAGNet (Yang et al. 2023b), with an improvement of +1.97 in AUC and +2.58 in aIOU. Notably, our method also demonstrates impressive affordance capabilities in unseen point cloud prediction, outperforming the second-best method, IAGNet by 1.22 in AUC and LASO by 0.05 in aIOU. Additionally, 3D segmentation and tracking methods such as PMF (Zhuang et al. 2021), XMF (Aiello et al. 2022), and MBDF (Tan et al. 2021) perform significantly worse compared to the affordance method.
@@ B.3 Loss Functions
Consistent with IAGNet (Yang et al. 2023b), the heatmap loss $\mathcal{L}_{HM}$ consists of a focal loss (Lin et al. 2017) and a dice loss (Milletari et al. 2016), which are used to supervise the predicted affordance heatmap and the corresponding ground truth. This can be formulated as follows:
@@ B.4 Evaluation Metrics
Here, we provide a detailed explanation of the evaluation metrics used in our study. Following previous research, we adopt AUC (Lobo et al. 2008), aIOU (Rahman and Wang 2016), SIM (Swain and Ballard 1991), and MAE (Willmott and Matsuura 2005) as our evaluation metrics:
@@
    $\text{IOU}=\frac{\text{Intersction Area}}{\text{Union Area}}.$ (21)